# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v2

@@ 3.2 Attack on vADDG
The idea behind this attack is to create a pseudo-random characteristic function and implement it homomorphically. To build such a function, we follow the following procedure. First, the adversarial server receives the inputs $\textsf{ct}{}_{s}:=[\mathbf{x}_{s}]_{2}$ for all $1\leq s\leq\gamma$ where each $\mathbf{x}_{s}\in\mathbb{Z}_{2}^{128}$ is a 128-bit string. We design a homomorphic characteristic function $\chi:\mathcal{C}^{128}\to\mathcal{C},\textsf{ct}{}_{s}\mapsto\overline{\textsf{ct}{}_{s,u}}$ as follows: First, Let an index $u\in\{1,\dots,128\}$ be given. Next, for each ciphertext string $\textsf{ct}{}_{s}$, evaluate
$\overline{\textsf{ct}{}_{s,u}}:={\sf CPPBS}_{(2,3)}(\textsf{ct}{}_{s,u}).$
Then, this characteristic function $\chi:\textsf{ct}{}_{s}\mapsto\overline{\textsf{ct}{}_{s,u}}$ is a homomorphic pseudo-random function which outputs $[1]_{3}$ with approximate probability $1/2$, otherwise $[0]_{3}$. With these $\gamma$ outputs $(\overline{\textsf{ct}{}_{s,u}})_{1\leq s\leq\gamma}$, the adversary can try to forge the PRF $F$ by adding $(\overline{\textsf{ct}{}_{s,u}})_{1\leq s\leq\gamma}$ to the output strings $\left(\textsf{Eval}_{F(\mathbf{k,-})}(\textsf{ct}{}_{s})\right)_{1\leq s\leq\gamma}$. For example, for ${\bf t}=(1,0,\dots,0)$, we can forge the PRF by evaluating
$\textsf{Eval}_{F({\bf k},-)}(\textsf{ct}{}_{s})+{\bf t}\cdot\chi(\textsf{ct}{}_{s}).$
Note that $\chi(\textsf{ct}{}_{s})$ is the same for the same inputs and so this forgery is not detected if $\chi({\textsf{ct}{}_{s}})$ is 0 for the all $\beta$ verification values. The success probability is given in Theorem 1.
@@
2:     $\mathbf{t}\leftarrow\mathbb{Z}_{3}^{82}$ for $\mathbf{t}\neq({0,0,\dots,0})$triangleright Select the trits to forge.
@@
4:         $\overline{\textsf{ct}{}_{s,u}}\leftarrow{\sf CPPBS}_{(2,3)}([x_{s,u_{1}}]_{2})$
5:         $[\mathbf{z}_{s}]_{3}\leftarrow\textsf{Eval}_{F(\bf k,-)}(\textsf{ct}{}_{s})$triangleright $[\mathbf{z}_{s}]_{3}=([z_{s,1}]_{3},\dots,[z_{s,82}]_{3})$.
@@
For a given parameter $(\alpha,\beta,\nu)$, the expected probability that the depth one circuit in Algorithm 1 successfully forges the output is
@@
The forgery succeeds when $\overline{\textsf{ct}{}_{s,u}}=[0]_{3}$ for all $s$ correspond to the $\beta$ verification values and at least one $\overline{\textsf{ct}{}_{s,u}}=[1]_{3}$ for $s$ corresponds to $\alpha$ message values. Let $p$ be the probability that $\overline{\textsf{ct}{}_{s,u}}$ became an encryption of 1. In the Algorithm 1, the expected value of $p$ over all input values is $1/2$. The attack success probability is obtained by subtracting the probability that all values become zero, $\left(1-p\right)^{\alpha+\beta}$, from the probability that all $s$ corresponding to $\beta$ become zero, $(1-p)^{\beta}$. Thus, the success probability is given by $Q(p):=\left(1-p\right)^{\beta}-\left(1-p\right)^{\alpha+\beta}.$ Substituting $p=1/2$, we have the desired result $\left(1-\left(1/2\right)^{\alpha}\right)\cdot\left(1/2\right)^{\beta}.$
@@
The first parameter has an overhead size ($(\alpha\nu+\beta)/\alpha$) of approximately ×$\times 6.0$ but requires a larger set of inputs, meanwhile the second parameter has an overhead size of approximately ×$\times 10.4$ but requires a smaller set of inputs.
@@ 4 Cryptanalysis on Replication Encoding
We recall the Replication Encoding (REP) and its verification procedure described in [CKP+24] which is implemented within BFV. For the convenience of presentation, we describe our attack assuming the evaluation of a polynomial $f:\mathbb{\mathbb{Z}}_{t}\to\mathbb{\mathbb{Z}}_{t}$ on a ciphertext encrypted using homomorphic encryption $\textsf{Enc}:\mathbb{Z}_{t}^{N}\rightarrow\mathcal{R}^{2}_{q}$ where $t=p^{r}$ is a power of a prime. However, it is straightforward to extend it to a multivariate polynomial circuit $f:\mathbb{Z}_{t}^{\mu}\rightarrow\mathbb{Z}_{t}^{\nu}$ for $\mu,\nu\in\mathbb{Z}^{+}$ where an element in $\mathbb{Z}_{t}^{\mu}$ is considered as an (extended) slot element.
@@ 4.1 Replication Encoding
Let $f$ be a circuit that the client requested to evaluate. The REP proceeds as follows.
@@
    The slots indexed by $S^{c}=\{1,\dots,n\}\setminus S$ are filled with a same message $m$, and the complement slots indexed by $S$ are filled with $n/2$ random values $v_{i}\overset{\$}{\leftarrow}\mathbb{Z}_{t}$.
We call the slots indexed by $S$ as verification slots, and the slots indexed by $S^{c}$ as computation slots. After evaluating $f$ to the ciphertext, its decryption contains $f(v_{i})$’s in the verification slots and $f(m)$ in the computation slots. For verification, it is assumed that the client knows $f(v_{i})$’s and their integrity in advance, either by outsourcing the computation of $f(v_{i})$’s without encryption to a third party using existing verifiable computation techniques or by securely computing them himself. Note that the verification values are not published, unlike the VOPRF scheme in [ADDG24]. The client verifies as follows:
@@ 4.2 Attacks on REP
Learning or guessing $S\subset\{1,\dots,n\}$ is not easy. Instead of recovering $S$, we take a similar approach to the attack on vADDG: To learn $S$ in encrypted state by computing $\textsf{Enc}(J_{S})$ for $J_{S}:=(j_{1},\dots,j_{n})$ where $j_{i}=1$ if $i\in S$ and $j_{i}=0$ otherwise. Once the adversary recovers $\textsf{Enc}(J_{S})$, the ciphertext can be easily forged by computing the following:
${}_{\textsf{forged}}:=\mathsf{Mult}(\textsf{Enc}(J_{S}),\textsf{Eval}_{f}(\textsf{ct}{}))+\mathsf{Mult}(\textsf{Enc}(J_{S^{c}}),\textsf{Eval}_{g}(\textsf{ct}{})),$ ct ) ) , (1)
where $f$ is the circuit that the client requested to evaluate, and $g$ is any modified circuit by the adversary. (Note that $\textsf{Enc}(J_{S^{c}})=\mathbf{1}^{N}-\textsf{Enc}(J_{S})$, thus computing $\textsf{Enc}(J_{S^{c}})$ is equivalent to computing $\textsf{Enc}(J_{S})$.)
Then $\textsf{ct}{}_{\textsf{forged}}$ provides the client with a malicious result $g(m)$ but successfully passes the verification step.
@@
  • $\textsf{Compare}(x,y):\mathbb{Z}_{t}^{2}\to\mathbb{Z}_{t}$ is $1$ if $x=y$ and 00 if $x\neq y$.
We note that $\textsf{Eval}_{\textsf{Compare}}$ in BFV can be computed by Fermat’s little theorem when its plaintext modulus $t$ is a prime, with $\lceil\log t\rceil$ homomorphic multiplications. But if $t=p^{r}$ for an integer $r>1$, we need a little modification.
Lemma 1.
$\mathsf{Eval}_{\mathsf{Compare}}$ can be executed within $O(\log t)$ homomorphic multiplications in BFV, both for $t=p$ and $t=p^{r}$ where $r>1$.
Proof.
Consider $e(x)=1-x^{\phi(t)}$ where $\phi$ is the Euler’s totient function so that $e(x-y)=0$ if $x-y\in\mathbb{Z}_{t}^{\times}$, and otherwise $e(x-y)=1$. If $t=p$, $e(x-y)$ is the desired $\mathsf{Compare}(x,y)$ function. Now suppose that $t=p^{r}$ and define a polynomial $\alpha$ as $\alpha(x):=e(x)\cdot\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u).$
Note that $\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u)=\frac{(x+1)^{\phi(t)}-1}{(x+1)-1}=\sum_{i=0}^{\phi(t)-1}(x+1)^{i},$ thus $\alpha$ can be evaluated within $O(\log t)$ homomorphic multiplications. Now we have three possible cases:
  1. 1.
    If $x=0$, then $\alpha(0)=e(0)\cdot\phi(t)=\phi(t)$.
  2. 2.
    If $x\in\mathbb{Z}_{t}^{\times}$, then $e(x)=0$ thus $\alpha(x)=0$.
  3. 3.
    If $x$ is a nonzero and non-unit, then $x+1\in\mathbb{Z}_{t}^{\times}\setminus\{1\}$. Thus $\alpha(x)=e(x)\left(\prod_{u\in\mathbb{Z}_{t}^{\times}\setminus\{1\}}(x+1-u)\right)=e(x)\cdot 0=0.$
Therefore for $x,y\in\mathbb{Z}_{t}$, interpolating $\alpha(x-y)$ as $\phi(t)\mapsto 1$ and $0\to 0$ would give the desired $\textsf{Compare}(x,y)$ function. ∎
Now we introduce Algorithm 2 to construct the desired circuit and describe its properties.
@@
1:Input: A ciphertext ct encrypted via REP and an index $i$.
@@
4:     ${}^{\prime}\leftarrow\mathsf{\textsf{Eval}_{Duplicate}}_{i}(\textsf{ct}{})$ ct ) $\triangleright$ $\textsf{ct}{}^{\prime}=\textsf{Enc}((x_{i},\dots,x_{i}))$
5:     ${}_{\textsf{CVS},i}\leftarrow\textsf{Eval}_{\mathsf{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ ct )
@@
For $i\overset{\$}{\leftarrow}\{1,\dots,n\}$, $\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ in the algorithm 2 outputs $\mathsf{Enc}(J_{S^{c}})$ with probability of $|S|/n=1/2$ in $O(\log t)$ homomorphic multiplications and rotations.
@@
Let ct be an encryption of a vector that has $m$ in computational slots $S^{c}$. If $i\in S^{c}$, then ${}^{\prime}\leftarrow\textsf{Eval}_{\mathsf{Duplicate}_{i}}(\textsf{ct}{})$ ct ) is an encryption of $(m,\dots,m)$. Consequently, $\mathsf{Eval_{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ returns $\textsf{Enc}(J_{S^{c}})$.
@@
If the attack fails, (i.e. $i\in S$) then the output of $\textsf{ct}{}_{\sf forged}$ of 1 where $\textsf{Enc}(J_{S^{c}})\leftarrow\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ is detected by the client. Thus this attack does not show that the REP is not sound in the presence of the covert adversary.
@@
Now we show how to increase the success probability of the attack by repeating Algorithm 2 for different $i$’s. For any fixed $i$, if we sum up the values in the $n$ slots of $\textsf{ct}{}_{\textsf{cvs},i}$, it outputs $|S|$ or $1$ depending on whether $i\in S$ or not. i.e. the server can homomorphically distinguish whether $i\in S$ or not. Using this principle, by repeating the algorithm 2, the adversary can construct a cheating circuit that deterministically outputs $\textsf{Enc}(J_{S^{c}})$. See Algorithm 3.
@@
Interpolate :$:\begin{cases}1\mapsto 0\\ n/2\mapsto 1\\ \text{otherwise anywhere.}\\ \end{cases}\textsf{Normalize}_{k}:\begin{cases}0\mapsto 0\\ 1,\dots,k\mapsto 1,\\ \text{otherwise anywhere}\\ \end{cases}$
Note that $\textsf{Eval}_{\textsf{RotSum}}$ can be evaluated within $\lceil\log n\rceil$ homomorphic rotations, and $\textsf{Eval}_{\textsf{Interpolate}},\textsf{Eval}_{\textsf{Normalize}_{k}}$ can be evaluated within $O(1)$ and $O(\log k)$ homomorphic multiplication respectively.
@@
1:Input: A ciphertext ct encrypted via REP and an index $i$.
@@
5:     $\textsf{ct}{}_{\textsf{rs}}\leftarrow\textsf{RotSum}(\textsf{ct}{}_{\textsf{cvs},i})$triangleright $\textsf{ct}{}_{\textsf{rs}}=\textsf{Enc}(1,\dots,1)$ or $\textsf{Enc}(n/2,\dots,n/2)$
6:     $\textsf{ct}{}_{\textsf{bool},i}\leftarrow\textsf{Eval}_{\mathsf{Interpolate}}(\textsf{ct}{}_{\textsf{rs}})$
7:     return $\textsf{ct}{}_{\textsf{bool},i}$triangleright $\textsf{ct}{}_{\textsf{bool},i}=\textsf{Enc}(0,\dots,0)$ or $\textsf{Enc}(1,\dots,1)$
@@
15:     $\textsf{ct}{}^{\prime}\leftarrow\sum_{i=1}^{k}\sf{Mult}(\textsf{ct}{}_{\textsf{bool},i},\textsf{ct}{}_{\textsf{cvs},i})$
@@
Since the slots in $\textsf{ct}{}_{\textsf{cvs},i}$ contain $n/2$ ones if $i\in S^{c}$, and a single one if $i\in S$, applying $\textsf{Eval}_{\textsf{RotSum}}$ yields either $\textsf{Enc}\left(\frac{n}{2},\dots,\frac{n}{2}\right)$ or $\textsf{Enc}(1,\dots,1).$ Consequently, by interpolation, the output $\textsf{ct}{}_{\text{bool},i}$ is a vector of Boolean values, either $(1,\dots,1)$ or $(0,\dots,0)$. Thus, for $i=1,\dots,k$, the sum of the $\textsf{ct}{}_{\textsf{cvs},i}$ after homomorphic multiplication with $\textsf{ct}{}_{\text{bool},i}$ decrypts to the same value as $|S^{c}\cap\{1,\dots,k\}|\cdot\textsf{Enc}(J_{S^{c}}).$ Since $S\subset\{1,\dots,n\}$ is a random subset of size $n/2$, $|S^{c}\cap\{1,\dots,k\}|$ is zero with probability ($\frac{{n/2\choose k}}{{n\choose k}}$ and otherwise nonzero. Thus, after evaluating $\textsf{Normalize}_{k}$, we obtain $\textsf{Enc}(0,\dots,0)$ with probability ($\frac{{n/2\choose k}}{{n\choose k}}$, or $\textsf{Enc}(J_{S^{c}})$ with probability $1-\frac{{n/2\choose k}}{{n\choose k}}$.
@@
By taking $k=n/2+1$, this attack can achieve a $100\%$ success rate [^2] With only negligible probability of failure, due to the properties of the underlying RLWE scheme; see Definitions 2 and 4. Even when $k<n$, the failure probability of the attack is bounded by ($\frac{{n/2\choose k}}{{n\choose k}}\leq 2^{-k}$, which decreases exponentially with $k$. Therefore, if the adversary is willing to trade off success probability for reduced circuit latency, they may choose a sufficiently small value of $k$. Moreover, since the algorithm ${\sf Recover}_{S^{c}}(\textsf{ct}{},k)$ is parallelizable, increasing $k$ may not lead to increased latency, depending on the adversary’s computational capabilities.

### v3

@@ 3.2 Attack on vADDG
The idea behind this attack is to create a pseudorandom characteristic function and implement it homomorphically. To build such a function, we follow the following procedure. Let the adversarial server receives the inputs $\textsf{ct}{}_{s}:=[\mathbf{x}_{s}]_{2}$ for all $1\leq s\leq\gamma$ where each $\mathbf{x}_{s}=(x_{s,u})_{1\leq u\leq 128}\in\mathbb{Z}_{2}^{128}$ is a 128-bit string. Now design a homomorphic characteristic function $\chi_{u}:\mathcal{C}^{128}\to\mathcal{C},\textsf{ct}{}_{s}\mapsto\overline{\textsf{ct}{}_{s,u}}$ as follows: First, let an index $u\in\{1,\dots,128\}$ be fixed. Next, for each ciphertext string $\textsf{ct}{}_{s}$, evaluate
[…]
Then, this characteristic function $\chi_{u}$ is a homomorphic pseudorandom function which outputs $1]_{3}$ with probability approximately $1/2$, otherwise $0]_{3}$. With these $\gamma$ outputs $(\overline{\textsf{ct}{}_{s,u}})_{1\leq s\leq\gamma}$, the adversary can try to forge the PRF $F$ by adding $(\overline{\textsf{ct}{}_{s,u}})_{1\leq s\leq\gamma}$ to the output strings $\left(\textsf{Eval}_{F(\mathbf{k,-})}(\textsf{ct}{}_{s})\right)_{1\leq s\leq\gamma}$. For example, for ${\bf t}=(1,0,\dots,0)$, we can forge the PRF by evaluating
[…]
@@
2:  $\mathbf{t}\leftarrow\mathbb{Z}_{3}^{82}$ for $\mathbf{t}\neq({0,0,\dots,0})$ $\triangleright$ Select the trits to forge.
@@
4:   $\overline{\textsf{ct}{}_{s,u}}\leftarrow{\sf CPPBS}_{(2,3)}([x_{s,u}]_{2})$
[…]
@@
For a given parameter $(\alpha,\beta,\nu)$, the expected probability that the bootstrapping-depth one circuit in Algorithm 1 successfully forges the output is
@@
The forgery succeeds when $\overline{\textsf{ct}{}_{s,u}}=[0]_{3}$ for all $s$ correspond to the $\beta$ verification values and at least one $\overline{\textsf{ct}{}_{s,u}}=[1]_{3}$ for $s$ corresponds to $\alpha$ message values. Let $p$ be the probability that $\overline{\textsf{ct}{}_{s,u}}$ became an encryption of 1. In Algorithm 1, the expected value of $p$ over all input values is $1/2$. The attack success probability is obtained by subtracting the probability that all values become zero, $\left(1-p\right)^{\alpha+\beta}$, from the probability that all $s$ corresponding to $\beta$ become zero, $(1-p)^{\beta}$. Thus, the success probability is given by $Q(p):=\left(1-p\right)^{\beta}-\left(1-p\right)^{\alpha+\beta}.$ Substituting $p=1/2$, we have the desired result $\left(1-\left(1/2\right)^{\alpha}\right)\cdot\left(1/2\right)^{\beta}.$
@@
The first parameter has an overhead size ($(\alpha\nu+\beta)/\alpha$) of approximately $\times 6.0$ but requires a larger set of inputs, meanwhile the second parameter has an overhead size of approximately $\times 10.4$ but requires a smaller set of inputs.
@@ 4 Cryptanalysis on Replication Encoding
We recall the Replication Encoding (REP) and its verification procedure described in [CKP+24]. For the convenience of presentation, we describe our attack assuming the evaluation of a single-input and single-output polynomial $f:\mathbb{Z}_{t}\to\mathbb{Z}_{t}$ on a ciphertext encrypted under the HE scheme $\textsf{Enc}:\mathbb{Z}_{t}^{N}\rightarrow\mathcal{R}_{q}^{2}$, where $t=p$ is a prime. However, it is straightforward to extend our attack to a multivariate polynomial circuit $f:\mathbb{Z}_{t}^{\mu}\rightarrow\mathbb{Z}_{t}^{\nu}$ for $\mu,\nu\in\mathbb{Z}^{+}$, where an element of $\mathbb{Z}_{t}^{\mu}$ is regarded as an extended slot element. For the case $t=p^{r}$ with $r>1$, a similar idea can be applied to construct an analogous attack. We defer the details for $t=p^{r}$ to Appendix 0.B.
@@ 4.1 Replication Encoding
Let $f:\mathbb{Z}_{t}\to\mathbb{Z}_{t}$ be a circuit that the client requested to evaluate. The REP proceeds as follows.
@@
    The slots indexed by $S^{c}=\{1,\dots,n\}\setminus S$ are filled with a same message $m$, and the complement slots indexed by $S$ are filled with $n/2$ random values $v_{i}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{t}$.
We call the slots indexed by $S$ as verification slots, and the slots indexed by $S^{c}$ as computation slots. After evaluating $f$ to the ciphertext, its decryption contains $f(v_{i})$’s in the verification slots and $f(m)$ in the computation slots. For verification, it is assumed that the client knows $f(v_{i})$’s and their integrity in advance, either by outsourcing the computation of $f(v_{i})$’s without encryption to a third party using existing verifiable computation techniques or by securely computing them himself. Note that the verification values are not published, unlike vADDG. The client verifies as follows:
@@ 4.2 Attacks on REP
Learning or guessing $S\subset\{1,\dots,n\}$ is not easy. Instead of recovering $S$, we take a similar approach to the attack on vADDG: we learn $S$ in encrypted state by computing $\textsf{Enc}(J_{S})$ for $J_{S}:=(j_{1},\dots,j_{n})$ where $j_{i}=1$ if $i\in S$ and $j_{i}=0$ otherwise. Once the adversary recovers $\textsf{Enc}(J_{S})$ [^2] Note that $J_{S^{c}}=(1,\dots,1)-J_{S}$, thus recovering $\textsf{Enc}(J_{S^{c}})$ is equivalent to recovering $\textsf{Enc}(J_{S})$., the ciphertext can be easily forged by computing the following:
$\textsf{ct}{}_{\textsf{forged}}:=\mathsf{Mult}(\textsf{Enc}(J_{S}),\textsf{Eval}_{f}(\textsf{ct}{}))+\mathsf{Mult}(\textsf{Enc}(J_{S^{c}}),\textsf{Eval}_{g}(\textsf{ct}{})),$ (1)
where $f$ is the circuit that the client requested to evaluate, and $g$ is any modified circuit by the adversary. Then $\textsf{ct}{}_{\textsf{forged}}$ provides the client with a malicious result $g(m)$ but successfully passes the verification step.
@@
  • $\textsf{Compare}(x,y):\mathbb{Z}_{t}^{2}\to\mathbb{Z}_{t}$ is $1$ if $x=y$ and 0 if $x\neq y$. We can extend it to act on vectors componentwise manner, $\textsf{Compare}:\mathbb{Z}_{t}^{2n}\to\mathbb{Z}_{t}^{n}$
Also we extend Compare to act on vectors componentwise manner. We note that Fermat’s little theorem guarantees $1-(x-y)^{t-1}=\textsf{Compare}(x,y)$ when $t$ is a prime, which needs $\lceil\log t\rceil$ multiplications to evaluate. Now we introduce Algorithm 2 to construct the desired circuit and describe its properties.
@@
1:Input: A ciphertext ct encrypted via REP and an index ii.
@@
4:  $\textsf{ct}{}^{\prime}\leftarrow\mathsf{\textsf{Eval}_{Duplicate}}_{i}(\textsf{ct}{})$ $\triangleright$ $\textsf{ct}{}^{\prime}=\textsf{Enc}((x_{i},\dots,x_{i}))$
5:  $\textsf{ct}{}_{\textsf{CVS},i}\leftarrow\textsf{Eval}_{\mathsf{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$
@@
For uniformly random $i\overset{\mathdollar}{\leftarrow}\{1,\dots,n\}$, $\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ in Algorithm 2 outputs $\mathsf{Enc}(J_{S^{c}})$ with probability of $|S|/n=1/2$ in $O(\log t)$ homomorphic multiplications and rotations.
@@
Let ct be an encryption of a vector that has $m$ in computational slots $S^{c}$. If $i\in S^{c}$, then $\textsf{ct}{}^{\prime}\leftarrow\textsf{Eval}_{\mathsf{Duplicate}_{i}}(\textsf{ct}{})$ is an encryption of $(m,\dots,m)$. Consequently, $\mathsf{Eval_{Compare}}(\textsf{ct}{}^{\prime},\textsf{ct}{})$ returns $\textsf{Enc}(J_{S^{c}})$.
@@
If the attack fails, (i.e. $i\in S$) then the output of $\textsf{ct}{}_{\sf forged}$ of the forgery fomular 1 where $\textsf{Enc}(J_{S^{c}})\leftarrow\mathsf{CVS}(\mathsf{\textsf{ct}{}},i)$ is detected by the client. Thus this attack does not imply that the REP is not sound in the presence of the covert adversary.
@@
Now we show how to increase the success probability of the attack by repeating Algorithm 2 for different indices $i$. For any fixed $i$, if we sum up the values in the $n$ slots of $\textsf{ct}{}_{\textsf{cvs},i}$, it outputs $|S|$ or $1$ depending on whether $i\in S$ or not. That is, the server can homomorphically distinguish whether $i\in S$ or not. Therefore, an adversary can construct a cheating circuit (Algorithm 3) that deterministically outputs $\textsf{Enc}(J_{S^{c}})$. To design such a circuit, first we define three functions $\textsf{RotSum},\textsf{Interpolate}$ and $\textsf{Normalize}_{k}$ as follows:
@@
Note that $\textsf{Eval}_{\textsf{RotSum}}$ can be evaluated using $\lceil\log n\rceil$ homomorphic rotations, while $\textsf{Eval}_{\textsf{Interpolate}}$ and $\textsf{Eval}_{\textsf{Normalize}_{k}}$ can be evaluated using $O(1)$ and $O(\log k)$ homomorphic multiplications, respectively, when the plaintext modulus $t$ is prime.
@@
1:Input: A ciphertext ct encrypted via REP and an index ii.
@@
[…]
7:  return $\textsf{ct}{}_{\textsf{bool},i}$ $\triangleright$ $\textsf{ct}{}_{\textsf{bool},i}=\textsf{Enc}(0,\dots,0)$ or $\textsf{Enc}(1,\dots,1)$
@@
[…]
@@
Since the slots in $\textsf{ct}{}_{\textsf{cvs},i}$ contain $n/2$ ones if $i\in S^{c}$, and a single one if $i\in S$, applying $\textsf{Eval}_{\textsf{RotSum}}$ yields either $\textsf{Enc}\left(\frac{n}{2},\dots,\frac{n}{2}\right)$ or $\textsf{Enc}(1,\dots,1).$ Consequently, by interpolation, the output $\textsf{ct}{}_{\text{bool},i}$ is a vector of Boolean values, either $(1,\dots,1)$ or $(0,\dots,0)$. Thus, for $i=1,\dots,k$, the sum of the $\textsf{ct}{}_{\textsf{cvs},i}$ after homomorphic multiplication with $\textsf{ct}{}_{\text{bool},i}$ decrypts to the same value as $|S^{c}\cap\{1,\dots,k\}|\cdot\textsf{Enc}(J_{S^{c}}).$ Since $S\subset\{1,\dots,n\}$ is a random subset of size $n/2$, $|S^{c}\cap\{1,\dots,k\}|$ is zero with probability $\frac{{n/2\choose k}}{{n\choose k}}$ and otherwise nonzero. Thus, after evaluating $\textsf{Normalize}_{k}$, we obtain $\textsf{Enc}(0,\dots,0)$ with probability $\frac{{n/2\choose k}}{{n\choose k}}$, or $\textsf{Enc}(J_{S^{c}})$ with probability $1-\frac{{n/2\choose k}}{{n\choose k}}$.
@@
By taking $k=n/2+1$, this attack can achieve a $100\%$ success rate [^3] With only negligible probability of failure, due to the properties of the underlying RLWE scheme; see Definitions 2 and 4.. Even when $k<n$, the failure probability of the attack is bounded by $\frac{{n/2\choose k}}{{n\choose k}}\leq 2^{-k}$, which decreases exponentially with $k$. Therefore, if the adversary is willing to trade off success probability for reduced circuit latency, they may choose a sufficiently small value of $k$. Moreover, since the circuit ${\sf Recover}_{S^{c}}(\textsf{ct}{},k)$ is parallelizable, increasing $k$ may not lead to increased latency, depending on the adversary’s computational capabilities.